OpenStudy (anonymous):

Solve D = 3(a – 2c) for a in terms of D and c.

hero (hero):

D = 3(a - 2c) D = 3a - 6c D - 6c = 3a (D - 6c)/3 = a

OpenStudy (anonymous):

D = 3(a - 2c) D = 3a - 6c D + 6c = 3a (D + 6c)/3 = a with all respect to Hero